Consider a Keynesian model:

Full employment output = R100 million

Tax rate = 0,25

Investment = R40 million

Autonomous consumption = R30 million

Marginal propensity to consume = 0,8

3.28 The value of the multiplier is …

[1] 2

[2] 1.67

[3] 2.5

[4] 4

3.29 The equilibrium level of income is …

[1] R70 million.

[2] R175 million.

[3] R280 million.

[4] R140 million.

3.30 To bring about full employment, government spending should be …

[1] -R30 million.

[2] -R72 million.

[3] R30 million.

[4] R75 million.


1
Expert's answer
2021-07-12T05:28:03-0400

3.28 The value of the multiplier is: m = \frac{1}{1 - 0.8(1 - 0.25)} = 2.5.

So, the correct answer is [3].

3.29 The equilibrium level of income is:

Y = C + I = 30 + 0.8(Y - 0.25Y) + 40 = 70 + 0.6Y,

0.4Y = 70,

Y = R175 million.

So, the correct answer is [2].

3.30 To bring about full employment, government spending should be:

(175 - 100)/2.5 = R30 million.

So, the correct answer is [3].